I have a 'wizard deck' of cards. Total of 79 cards (44 black and 35 red). You draw a two card hand. What's the probability that you will draw AT MOST one red card. I found the probability of getting one red card = 35/79 and the probability of getting one black = 44/78. And I multiplied so 1540/6162??? This is how I interpreted the question. Am I correct?
Answer by richard1234(7193)

It is equal to one minus the probability of drawing a hand with two red cards. This is because if you do not draw a red two-card hand, then you must have drawn a hand with 0 or 1 red cards.

Here, P(2 red cards) = (35/79)(34/78) = 595/3081. This implies P(0 or 1 red cards) = 1 - 595/3081 = 2486/3081.